OpenStudy (secret-ninja):

Am I right? In the function f(x) = 5(x^2 − 4x + ____) + 15, what number belongs in the blank to complete the square?

OpenStudy (secret-ninja):

The number 4 belongs in the space because you get half of the coefficient and then square it, correct?

OpenStudy (secret-ninja):

I'm GREEEEEEN OMG. :D :D :D AS OF 12:22 PM OF 1/7/2015 AAAHHHHHH

OpenStudy (secret-ninja):

LOL

OpenStudy (secret-ninja):

@ganeshie8 @undeadknight26 @blurbendy

OpenStudy (mathmath333):

\(\Huge \checkmark\)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{(x+2)^2=x^2+4x+4\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (secret-ninja):

Yay thanks ^-^

OpenStudy (mathmath333):

sry \(\large \begin{align} \color{black}{(x-2)^2=x^2-4x+4\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (secret-ninja):

Can you help me in one more? I'll tag you.